OpenStudy (anonymous):

A line has a slope of 0.5 and a y-intercept of (0, -2). What is the value of y when x = 4?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

@ikram002p

OpenStudy (ikram002p):

find line equation (y--2)=m(x-0) s.t m is the slope y+2=0.5x y=0.5x-2 when x=4 y=0.5(4)-2=2-2=0

OpenStudy (anonymous):

so the answer is 0?

OpenStudy (ikram002p):

yess

OpenStudy (anonymous):

thank you:)

OpenStudy (ikram002p):

\(\Huge\color{red}{ \text{np ت}}\)

ganeshie8 (ganeshie8):

lol ur boasting ur latex skills ha ;p

OpenStudy (ikram002p):

heh XD